# Prioritizing Security Practice Adoption: Empirical Insights on Software Security Outcomes in the npm Ecosystem

### v1

@@ Abstract.
Practitioners often struggle with the overwhelming number of security practices outlined in cybersecurity frameworks for risk mitigation. Given the limited budget, time, and resources, practitioners want to prioritize the adoption of security practices based on empirical evidence. The goal of this study is to assist practitioners and policymakers in making informed decisions on which security practices to adopt by evaluating the relationship between software security practices and security outcome metrics. The study investigated the relationship between security practice adoption and security outcomes. We selected the OpenSSF Scorecard metrics to automatically measure the adoption of security practices in npm GitHub repositories. We also explored security outcomes metrics, such as the number of open vulnerabilities (Vul_Count), mean time to remediate (MTTR) vulnerabilities in dependencies, and mean time to update (MTTU) dependencies. We conducted regression and causal analysis using 12 Scorecard metrics and their aggregated Scorecard score (computed by aggregating individual security practice scores) as predictors and Vul_Count, MTTR, and MTTU as target variables. Our initial dataset included 146K npm packages that have Scorecard, Vul_Count , and MTTU data. However, 85% lacked MTTR data because their dependencies either had no reported vulnerabilities or had vulnerabilities that remained unpatched. To address this, we extracted a 22K subset from the 146K dataset with MTTR data. Our findings show that higher aggregated Scorecard scores are associated with fewer Vul_Count and shorter MTTU, also supported by causal analysis. However, while the regression model suggests shorter MTTR, causal analysis indicates project characteristics likely influence MTTR direction. Segment analysis shows that larger, newer repositories with more contributors, dependencies, and downloads have shorter MTTR. Among individual security practices, Code Review, Maintained status, Pinned Dependencies, and Branch Protection show strong associations with security outcomes; the directionality of these associations varies across security outcomes. Additionally, MTTU follows trends similar to Vul_Count and MTTR, making it a potential proxy metric when vulnerability and MTTR data are unavailable.
@@ 1. Introduction
The package release frequency has increased by 1,466% over the last decade, with smaller applications now incorporating an average of 180 components (sonatype2024, ). The growing reliance on third-party dependencies in modern applications expands the attack surface, exposing applications to CVE (Common Vulnerabilities and Exposures catalog) exploits or malicious code injection, facilitating software supply chain (SSC) attacks (williams2024research, ). For example, CVEs have grown by 463% in the last 10 years, and 704,102 malicious packages have been proactively identified since 2019 (sonatype2024, ).
Practitioners are falling behind in managing vulnerabilities in dependencies due to growing dependency complexity and resource constraints, with mean time to remediate (MTTR) increasing from under 25 days in 2017 to over 400 days by 2024 (sonatype2024, ). SSC security standards have been proposed by industry and government agencies to help organizations adopt security practices that mitigate SSC security risks. Security practices are actions, procedures, techniques, or other measures that an organization can adopt to reduce the vulnerability of an information system (NISTSecurityControls, ). Examples of security practices include conducting code reviews before merging pull requests and using a static analyzer tool to detect vulnerabilities. In this paper, we used the term “security practice,” also synonymous with security control, safeguards, and countermeasures (NISTSecurityControls, ).
Many stakeholders have recognized SSC security concerns; most prominently, the Executive Order (EO) 14028 (EO_2021, ) mandates the adoption and self-attestation of security practices to improve SSC security. Similarly, the European Union’s Cyber Resilience Act (CRA), approved in March 2024, mandates common cybersecurity standards for digital products, including incident reporting and automatic MTTR updates (ENISA_CRA_2024, ). In 2025, the US EO 14144 (EO2025Cybersecurity, ) strengthens the efforts by requiring- “Within 240 days of the date of this order,….the Director of NIST, shall evaluate common cybersecurity practices and security outcomes that are commonly used or recommended across industry sectors.” Security outcomes are the indicators of the performance or non-performance of security practices prescribed for an information system (EO2025Cybersecurity, ).
While the SSC security standards (EO_2021, ; ENISA_CRA_2024, ; EO2025Cybersecurity, ; souppaya2022secure, ; Microsoft_framework, ; Williams2024PSSCRM, ) contain a long list of security practices that practitioners should adopt, practitioners are often overwhelmed by the number of practices recommended in these standards (zahan2023software, ). Given the limited budget, time, and resources (sammak2023developers, ; hamer2025closing, ), practitioners want to prioritize adopting security practices based on empirical evidence showing that adopting security practices mitigates security risk. The goal of this study is to assist practitioners and policymakers in making informed decisions on which security practices to adopt by evaluating the relationship between software security practices and security outcome metrics. Our research directly aligns with federal mandates EO 14144. However, instead of identifying commonly adopted security practices, we investigate the relationship between security practices and outcomes, providing empirical evidence to understand which practices are associated with improved security. We address the following research questions (RQs) and hypotheses:
@@
  • RQ1: Is higher adoption of security practices associated with measurable improvements in security outcomes?
@@
To that end, the study investigated the npm ecosystem to understand the relationship between publicly available data on security practice adoption and security outcomes. We selected the OpenSSF Scorecard (Scorecard, ) tool for our large-scale analysis because the tool utilizes metrics that automatically measure the adoption of security practices in the GitHub repositories. The Scorecard (Scorecard, ) uses 18 predefined metrics to detect the adoption of security practices. For each metric, the Scorecard tool assigns a score of 0-10 based on pre-defined criteria. The tool also computes an aggregate security score as a weighted average of the 18 individual security metrics to understand the overall security of the repositories. For security outcomes, we selected the number of open vulnerabilities (Vul_Count), the mean time to remediate (MTTR), and the mean time to update (MTTU) metrics. To address RQ1, we developed a regression model to assess the relationship between a package’s aggregate security score and security outcome metrics. Additionally, we applied Propensity Score Matching (PSM), a statistical matching technique used in observational studies, to estimate causal effects and mitigate the potential confounding effects. For RQ2, we employed a regression model and feature importance-ranked techniques to identify which security practices have the strongest associations with security outcomes. To further validate these findings, we used PSM to estimate the effect of individual security practices on security outcomes, isolating their impact while accounting for differences in repository characteristics.
Our research builds on Zahan et al. (zahan2023software, ), which investigated whether OpenSSF Scorecard metrics predict vulnerabilities in the PyPI and npm ecosystems. Their study found a counterintuitive positive association between security practices and vulnerabilities and attributed the result to a small dataset (2,422 packages). However, the authors collected vulnerabilities at the package level while obtaining security practices from the GitHub repositories, which often contain additional dependencies and languages, leading to incomplete vulnerability data. We improved data collection by 1) incorporating vulnerabilities from dependencies and repositories instead of package level; 2) constructing a large dataset (146k npm packages); 3) including control variables (project age, size, etc.) to address external factors that might have an influence; and 4) introducing MTTR and MTTU as novel security outcome metrics to provide a more comprehensive evaluation. The study will have the following contributions:
  • MTTR and MTTU metrics to measure security outcomes.
@@
  • Empirical evidence on how security practices impact security outcome of npm packages.
The paper is organized as follows: Section 2 discusses metrics used in our study, and Section 3 discusses our dataset construction steps. Section 4, Section 5 and Section 6 describes the data processing, methodology and result. We close with a discussion of our study (Section 7) and related work in Section 8.
2. Metrics Utilized in the Study
@@
2.1. Security Practice Metrics
In this study, we conducted a large-scale analysis to understand the adoption of security practices in OSS. We used the OpenSSF Scorecard tool (Scorecard, ), which evaluates security practices that can be measured automatically in GitHub. The tool detects security practice adoption, such as whether the package has implemented code review, repositories have enabled branch protection, or the repo is actively maintained. Table  1 contains the list of 18 security practices measured by Scorecard. Beyond industry adoption, the primary academic research on OpenSSF Scorecard has been conducted in study (zahan2023openssf, ; zahan2023software, ). The authors examine the adoption trends of security practices across ecosystems as measured by Scorecard metrics, analyze the distribution of Scorecard scores, and evaluate the efficacy of the Scorecard tool in assessing security practices. The author also verified whether the practices recommended by the Scorecard tool are also recommended in the standard security framework SSDF provided by NIST. The P-SSCRM SSC Framework (Williams2024PSSCRM, ) also included the OpenSSF Scorecard as one of the industry standards.
2.2. Security Outcome Metrics
Security outcomes are measured through quantitative characteristics, providing insights into the effectiveness of security practice adoption (scottmadden_outcome_metrics, ). In this study, we selected two temporal vulnerability metrics that capture both exposures (Vul_Count: number of open vulnerabilities) and responsiveness (MTTR: mean time to remediate vulnerable dependencies) and one temporal software metric (MTTU: mean time to update dependencies) as security outcome metrics.
  • Vul_Count: A measure quantifying the number of publicly disclosed or reported open security vulnerabilities within a project and its dependencies.
@@ 2.2. Security Outcome Metrics
The number of vulnerabilities is widely used in academic research, industry, and policy as outcome metrics (EO2025Cybersecurity, ; CISASecureSoftwareAttestation, ; morrison2014mapping, ; sec_outcome, ; morrison2015security, ; zahan2023software, ). MTTR is most commonly used to measure the package’s security (MTTR_1, ; MTTR_2, ), measuring the code dependencies security (MTTR_3, ) in the industry. The OpenSSF Secure Supply Chain Consumption Framework (S2C2F) framework (Microsoft_framework, ) includes MTTR as a key maturity indicator, while Sonatype (sonatype-2019, ) has used MTTR to classify secure OSS projects. However, measuring MTTR poses a challenge since many package dependencies either have no reported vulnerabilities or remain unpatched, limiting the dataset size. To mitigate this limitation, we also consider MTTU to be a proxy metric for MTTR. Rahman et al. (rahman2024characterizing, ) found a moderate correlation between MTTR and MTTU, suggesting that MTTU can used as a proxy when MTTR data is unavailable. We consider MTTU as an alternative indicator of MTTR: 1) MTTU data was derivable for all components; and 2) We hypothesize developers staying up to date on dependencies will also stay up-to-date on security updates because security updates are a subset of regular updates. Hence, packages will exhibit similar MTTR and MTTU values. The hypothesis, data availability, and the observed correlation between MTTR and MTTU (rahman2024characterizing, ) motivated us to consider MTTU as an additional security outcome metric.
2.3. Control Variable Metrics
Control variables are properties that researchers hold constant for all observations in an experiment to understand the true relationships between the predictor and target variables (FrostControlVariables, ). In this research, we evaluated the impact of security practices on security outcomes while including other factors that may also influence the outcome. We included five control variables to measure project characteristics. GitHub contributions count controls for developer activity, while package downloads account for package popularity. Repository age (years) controls the project’s development stage, and size (KB) controls the complexity of maintaining a larger package. Direct dependency count controls a project’s reliance on and exposure to external dependencies.
@@ 2.3. Control Variable Metrics
Rows with bolded security practice names indicate excluded metrics, either due to high computational expense (HCE) or because more than 90% of repositories had a score of -1. A score of -1 indicates that Scorecard could not get conclusive evidence of implementing practices, or perhaps an internal error occurred due to a runtime error. A score between 0-10 represents the extent of adoption of the security practice.
Security Practices Description Risk Label -1 score 0-10 score
Dangerous-Workflow Does the project avoid dangerous coding patterns in GitHub Action workflows? Critical 73.2% 26.8%
Branch-Protection Does the project use Branch Protection? High ¡1% (7) 99.9%
Code-Review Does the project practice code review before code is merged? High ¡1.4% (2071) 98.6%
Binary-Artifacts Is the project free of checked-in binaries? High ¡1% (3) 99.9%
@@
Dependency-Update-Tool Does the project use tools to help update its dependencies? High HCE HCE
Maintained Is the project at least 90 days old and maintained? High 0 100%
Signed-Releases Does the project cryptographically sign releases? High 97.7% 2.3%
Token-Permissions Does the project declare GitHub workflow tokens as read only? High 74.4% 26.8%
Vulnerabilities Does the project have unfixed vulnerabilities (including depedencies)? High 0 100%
Pinned-Dependencies Does the project declare and pin dependencies? Medium 72.2% 27.8%
Packaging Does the project build and publish official packages from CI/CD, e.g. GitHub Publishing ? Medium 97.4% 2.5%
Fuzzing Does the project use fuzzing tools, e.g. OSS-Fuzz, QuickCheck or fast-check? Medium ¡1% (15) 99.9%
@@
Static Application Security Testing (SAST) Does the project use static code analysis tools, e.g. CodeQL, LGTM (deprecated), SonarCloud? Medium ¡1% (203) 99.9%
Security-Policy Does the project contain a security policy? Medium ¡1% (3) 99.9%
CI-Tests Does the project run tests in CI, e.g. GitHub Actions, Prow? Low HCE HCE
Contributors Does the project have contributors from at least two different organizations? Low HCE HCE
License Does the project declare a license? Low ¡1% (2) 99.9%
CII-Best-Practices Has the project earned an OpenSSF (formerly CII) Best Practices Badge at the passing, silver, or gold level? Low ¡1% (12) 99.9%
3. Dataset Construction
The primary source of our dataset are the npm packages and their dependency graph and GitHub repositories. In Section 3.1, we discuss different sources used to construct our dataset. The package inclusion and exclusion criteria are discussed in Section 3.2. Section  3.3 discusses our security practice metrics data collection process using the OpenSSF Scorecard tool. Section 3.4 and Section 3.5 cover our security outcome metrics and control variables.
3.1. Data Source

### v2

# Assumptions to Evidence: Evaluating Security Practices Adoption and Their Impact on Outcomes in the npm Ecosystem

@@ Abstract.
Practitioners often struggle with the overwhelming number of security practices outlined in cybersecurity frameworks for risk mitigation. Given the limited budget, time, and resources, practitioners want to prioritize the adoption of security practices based on empirical evidence. The goal of this study is to assist practitioners and policymakers in making informed decisions on which security practices to adopt by evaluating the relationship between software security practices adoption and security outcome metrics. To do this, we analyzed the adoption of security practices and their impact on security outcome metrics across 145K npm packages. We selected the OpenSSF Scorecard metrics to automatically measure the adoption of security practices in npm GitHub repositories. We also investigated project-level security outcome metrics: the number of open vulnerabilities (Vul_Count), mean time to remediate (MTTR) vulnerabilities in dependencies, and mean time to update (MTTU) dependencies. We conducted regression and causal analysis using 11 Scorecard metrics and the aggregated Scorecard score (computed by aggregating individual security practice scores) as predictors and Vul_Count, MTTR, and MTTU as target variables. Our findings reveal that aggregated adoption of security practices is associated with 5.2 fewer vulnerabilities, 216.8 days faster MTTR, and 52.3 days faster MTTU. Repository characteristics have an impact on security practice effectiveness: repositories with high security practices adoption, especially those that are mature, actively maintained, large in size, have many contributors, few dependencies, and high download volumes, tend to exhibit better outcomes compared to smaller or inactive repositories. Among individual security practices, Code Review, Contributors from diverse organizations, License, CI-Tests, and Pinned Dependencies show strong associations with security outcomes.
@@ 1. Introduction
The growing reliance on third-party dependencies in modern software applications expands the attack surface, exposing applications to CVE (Common Vulnerabilities and Exposures catalog) exploits or malicious code injection, facilitating software supply chain (SSC) attacks (williams2024research, ). For example, CVEs have grown by 463% in the last 10 years, and 704,102 malicious packages have been proactively identified since 2019 (sonatype2024, ). Practitioners are falling behind in managing vulnerabilities in dependencies due to growing dependency complexity and resource constraints, with mean time to remediate (MTTR) vulnerabilities increasing from under 25 days in 2017 to over 400 days by 2024 (sonatype2024, ).
SSC security standards and frameworks (e.g.  (EO_2021, ; ENISA_CRA_2024, ; EO2025Cybersecurity, ; souppaya2022secure, ; Microsoft_framework, ; Williams2024PSSCRM, )) have been proposed by industry and government agencies to help organizations adopt security practices that mitigate SSC security risks. However, historical experience demonstrates the risks of adopting security practices without empirical validation. For example, long password complexity requirements, such as mandating uppercase letters, numbers, and symbols, were long assumed to enhance security (adams1999users, ; morris1979password, ). However, empirical studies (dekoven2022measuring, ; dekoven2019measuring, ; lain2024content, ) later revealed that such policies often led users to create predictable passwords by encouraging reuse and minor modifications. Similar theoretical assumptions may persist in SSC security, where different practices are widely promoted without comprehensive empirical validation. Security success is often measured through adoption metrics—e.g., the number of teams that adopted code review or the number of vulnerabilities detected or fixed, rather than actual security outcomes. Therefore, little is known about the effectiveness of best practices in improving security. The absence of empirical evidence arises from multiple systemic challenges: widespread disagreement about best practices among experts (reeder2017152, ; redmiles2020comprehensive, ), poor evidence even where consensus exists (shostack2019empirical, ; thompson2017large, ), and dependence on intuition and convention rather than empirical evidence.
The cybersecurity community’s reliance on perceived common sense has created an information-poor environment characterized by conflicting guidance and expert disagreement about fundamental priorities (reeder2017152, ; redmiles2020comprehensive, ; national2025cyber, ). Meanwhile, organizations have limited budgets, time, and resources (sammak2023developers, ; hamer2025closing, ), yet practitioners currently lack the empirical evidence needed to prioritize security practices that mitigate security risk. Therefore, developing empirical research to evaluate the security outcome of recommended practices is both a research priority and a fundamental prerequisite for evidence-based security decision-making (scott2025counting, ; national2025cyber, ). The goal of this study is to assist practitioners and policymakers in making informed decisions on which security practices to adopt by evaluating the relationship between software security practices adoption and security outcome metrics. We address the following research questions (RQs):
@@
  • RQ1: Is higher adoption of aggregated security practices associated with measurable improvements in security outcomes?
@@
To that end, the study investigated the relationship between publicly available data on security practice adoption and security outcomes metrics. For large-scale analysis, we focus on security practice adoption that can be measured automatically. We selected the OpenSSF Scorecard tool (Scorecard, ) that automatically collects 18 metrics on the adoption of security practices in GitHub repositories. The Scorecard metrics are valuable for collecting ecosystem-scale data, as well as for evaluating the security posture of project dependencies. Examples of security practices that can be measured by Scorecard include maintainers conducting code reviews before merging pull requests, using a static analyzer tool, and others. For each metric, the Scorecard tool assigns a score of 0-10 based on pre-defined criteria. The tool also computes an aggregate security score as a risk-weighted average of the individual security practice metrics, weighted by risk. For security outcomes, we selected project-based outcome metrics, where a project is defined as a software package that includes all components in its dependency graph. Specifically, the number of open vulnerabilities (Vul_Count), the mean time to remediate (MTTR) vulnerable and outdated dependencies, and the mean time to update (MTTU) outdated dependencies were used.
To address RQ1, we investigate the aggregate adoption of security practices and their impact on security outcomes. We developed a regression model to assess the relationship using Scorecard aggregate security score and security outcome metrics. Additionally, we applied Propensity Score Matching (PSM), a statistical matching technique used in observational studies, to estimate causal effects and mitigate the potential confounding effects based on repository characteristics. For RQ2, we employed a regression model and feature importance-ranked techniques to identify which individual security practices have the strongest associations with security outcomes. To further validate these findings, we used PSM to estimate the effect of individual security practices on security outcomes, isolating their impact while accounting for differences in repository characteristics. The study has the following contributions:
  • Introduction of project-based security outcome metrics.
  • Application of industry-standard metrics, MTTR and MTTU, to assess security outcomes.
@@
  • Empirical evidence on how security practices impact the security outcome of npm packages.
The paper is organized as follows: Section 2 contains the definition of security practice and security outcome; Section 3 discusses the design of our study; Section 3.1 discusses the metrics used in our study; and Section 3.2 discusses our dataset construction steps. Section 4, Section 5 and Section 6 describe the data processing, methodology, and results. We close with a discussion of our study (Section 7) and related work in Section 8.
2. Definition
Security practices are actions, procedures, and technical controls that an organization can adopt with the goal of preventing, detecting, and responding to threats against information systems. Security practices encompass both technical and organizational measures that are typically standardized and carried out by individuals or teams as part of an organization’s efforts to protect its information, resources, and operations (king2000best, ).
Security outcomes are quantifiable indicators that reflect the effectiveness of security practices implemented within an information system (EO2025Cybersecurity, ). Outcome metrics assess whether implemented security practices are achieving their intended security risk management objectives, specifically reducing harm to information systems and guiding continuous security improvement. Security outcome metrics may include indicators such as reductions in attack frequency, security incidents, or financial losses; decreased vulnerability exposure (e.g., lower vulnerability counts), and faster response times, which shorten the window of exposure to threats.
3. Experiment design
In this section, we outline the experimental design and dataset used to investigate our research questions. To answer RQ1 and RQ2, we define Scorecard security practices metrics and aggregate security score as predictor variables and security outcome metrics as target variables. We also incorporate control variables to account for potential confounding project characteristics. We leveraged machine learning and causal analysis techniques to examine the relationship between practices and outcomes. Our analysis is based on data collected from npm packages and their corresponding GitHub repositories, using dependency graphs to define project scope. The dataset construction integrates multiple data sources and applies systematic inclusion and exclusion criteria to ensure reliability and reproducibility. Section 3.1 details the selection of metrics used in our study, while Section 3.2 describes how we collected the dataset.
3.1. Metrics Utilized in the Study
@@
3.1.1. Security Practice Metrics
To assess the adoption of security practices, the OpenSSF Scorecard tool (Scorecard, ) was employed. The tool automatically evaluates a set of security practices implemented in GitHub repositories, based on observable project metadata and activity. Table  1 contains the list of 18 security practices measured by Scorecard. The tool also computes an aggregate security score to assess the overall security of a repository, computed as a risk-weighted average of individual security practice metrics. Scorecard was selected due to its widespread adoption and recognition in both academic and industry contexts. In academic research, Scorecard has been the subject of empirical evaluations, including SSC risk measurement (zahan2023software, ; mounesan2023exploring, ; butler2024links, ; salvievaluating, ; siddiqui2024elevating, ; kancharoendee2025categorizing, ; wu2025libvulnwatch, ; rydin2025evaluating, ), surveys on practitioners’ perspectives on Scorecard (elder2024applying, ; rydin2025evaluating, ), and qualitative studies (zahan2023openssf, ) examining its applicability and effectiveness. In industry, Scorecard has been adopted as part of standard security frameworks and requirements (Williams2024PSSCRM, ; openssf_osps_baseline_2025, ) and has also been used in reports by organizations, such as Sonatype and Veracode, to assess the state of open-source software ecosystems (sonatype_2022, ; eng2024quantifying, ). Adoption by major projects such as Kubernetes, Node.js, and Apache, along with integration into platforms like Ortelius (ortelius2024dashboard, ) and Deps.dev (OSI, ), also highlights wide adoption.
3.1.2. Security Outcome Metrics
A comprehensive understanding of security outcome metrics remains an open research challenge. Although defining and validating such metrics is inherently challenging, our work represents a step toward advancing that goal. In this study, we selected project-level outcome metrics because prior studies (zahan2023software, ) showed that only a small fraction of packages (0.1% of the npm ecosystem) have reported vulnerabilities at the individual package level. Limited availability of vulnerability data prevents models from learning meaningful relationships between security practices and security outcomes. Therefore, we leveraged three project-level outcome metrics, where a project is defined as an individual package and its dependencies. The security risk of a project decreases when a vulnerable component is replaced with a fixed one, providing a measurable indication of improved security outcomes, reflected in fewer open vulnerabilities and reduced exposure windows due to faster response times.
  • Vul_Count: A measure quantifying the number of publicly disclosed or reported open security vulnerabilities within a package and its full dependency tree, including both direct and transitive dependencies.
@@ 3.1.2. Security Outcome Metrics
The number of vulnerabilities is widely used in academic research, industry, and policy as outcome metrics (EO2025Cybersecurity, ; CISASecureSoftwareAttestation, ; morrison2014mapping, ; sec_outcome, ; morrison2015security, ; zahan2023software, ). While MTTR and MTTU are widely used industry metrics, our work is novel in applying them as outcome measures for security responsiveness. Faster updates reduce exposure windows, limiting attackers’ opportunity to exploit disclosed weaknesses. We draw inspiration from the DevOps Research and Assessment (DORA) framework (dora_metrics, ). DORA is one of the largest research studies in software engineering research, analyzing survey responses from over 31,000 professionals across six years (researched in the years 2014 through 2019) (forsgren2019accelerate, ). DORA research identifies which software engineering practices, including security practices, improve software delivery performance. One of DORA’s outcome metrics is a responsiveness metric that measures how long it takes a project to respond and deploy the corresponding updates. Research conducted in 2022 by the Google DORA team (google2022accelerate, ) specifically investigated SSC practices derived from the Secure Software Development Framework (SSDF) (souppaya2022secure, ) and Supply-chain Levels for Software Artifacts (SLSA) (SLSA, ), and examined their impact on outcomes. The research has shown that SSC practices are positively associated with faster response and deployment times, and practitioners also anticipate a lower chance of security vulnerabilities. Motivated by DORA metrics, we hypothesize that teams that adopt good security practices are also more likely to be more responsive, as measured by MTTR and MTTU. We used MTTU because ecosystem-wide MTTR data is often unavailable (only 15% of our dataset; see section 3.2.4). While MTTR focuses on the response to security updates, MTTU measures responsiveness to any upstream updates and is available for all packages. Prior work (rahman2024characterizing, ) also found a moderate correlation between MTTR and MTTU, suggesting that MTTU can serve as a proxy when MTTR is unavailable.
3.1.3. Control Variables
Control variables are properties that researchers hold constant in an experiment to understand the true relationships between the predictor and target variables (FrostControlVariables, ). We included six variables to measure repository characteristics based on two criteria: (1) alignment with prior use of control variables on software quality and practices work (miller2025understanding, ; zhou2019fork, ; ray2014large, ; ortu2019empirical, ; trockman2018adding, ; li2024can, ; javan2023dependency, ), and (2) maintaining model interpretability with a minimal yet effective set of metrics to reduce data collection and cost. We controlled for six repository characteristic metrics grouped by their hypothesized influence. Contributor count (Contributors_CT) and commit staleness reflect team capacity and development activity, which may affect the implementation and maintenance of security practices. Download count and direct dependencies capture project adoption and external exposure, potentially influencing both scrutiny and security risk. Repository age and size represent project maturity and complexity, which may impact maintenance effort and security performance. We excluded highly correlated variables such as forks, stars, and dependents due to redundancy and because they did not improve model performance.
@@ 3.1.3. Control Variables
-1: no conclusive evidence or runtime error. 0–10: extent of security practice adoption. Values in parentheses indicate raw counts.
Security Practices Description -1 % 0% 1–10%
Binary-Artifacts Is the project free of checked-in binaries? 0.0 0.2 99.8
Branch-Protection Does the project use Branch Protection? 0.0 (3) 85.9 14.1
Code-Review Does the project practice code review before code is merged? 0.4 75.4 24.2
License Does the project declare a license? 0.0 30.9 69.1
Maintained Is the project at least 90 days old and maintained? 0.0 86.7 13.3
@@
Dependency-Update-Tool Does the project use tools to help update its dependencies? 0.0 (65) 81.3 18.7
Contributors Does the project have contributors from different organizations? 0.0 36.7 63.3
@@
SAST Does the project use static code analysis tools, e.g. CodeQL? 0.0 (23) 96.9 3.1
Security-Policy Does the project contain a security policy? 0.0 (2) 93.2 6.8
CI-Tests Does the project run tests in CI, e.g. GitHub Actions, Prow? 50.3 41.3 8.4
Pinned-Dependencies Does the project declare and pin dependencies? 72.2 18.0 9.8
Token-Permissions Does the project declare GitHub workflow tokens as read-only? 73.2 25.6 1.2
Dangerous-Workflow Does the project avoid dangerous patterns in GitHub workflows? 73.2 0.3 26.6
Fuzzing Does the project use fuzzing tools, e.g. OSS-Fuzz, QuickCheck? 0.0 (2) 99.8 0.2
CII-Best-Practices Has the project earned an OpenSSF Best Practices Badge? 0.0 99.9 0.1
Signed-Releases Does the project cryptographically sign releases? 97.7 2.3 0.0
Packaging Does the project build and publish official packages from CI/CD? 97.5 0.0 2.5
Vulnerabilities Does the project have unfixed vulnerabilities (including dependencies)? 0.0 32.2 67.8
3.2. Dataset Source and Construction
Section 3.2.1 discusses different sources used to construct our dataset. Section 3.2.2, Section 3.2.3, Section 3.2.4 and Section 3.2.5 cover our dataset collection process. Our dataset is available in Figshare [^1] https://figshare.com/s/e74f94dfe1be6a90c4d1?file=54861686.
3.2.1. Data Sources